OpenStudy (anonymous):

please help me...An object is projected with an initial speed,v at an angle b, to the horizontal .Derive an equation for the path of the object.

OpenStudy (anonymous):



OpenStudy (anonymous):

Path for X axis: \[x_f=x_i+v \times \cos(b) \times t\] Path for Y axis: \[y_f=y_i+v \times \sin(b) \times t - \frac{ a \times t^2 }{ 2 }\]